OpenStudy (anonymous):

Louise has only 5 cent, 10 cent, and 20 cent coins in her purse. She has 30 coins in total and she has two more 10 cent coins than 5 cent coins. If the total value of her coins is $3.80, how many 10 cent coins does she have?

OpenStudy (aum):

Let the number of 5 cent coins be x. The number of 10 cent coins = x + 2 (because "she has two more 10 cent coins than 5 cent coins.") Number of 20 cent coins = 30 - x - (x+2) (because "She has 30 coins in total") Number of 20 cent coins = 30 - 2x - 2 = 28 - 2x Find the total value for each set of coins. Value in dollars of "x" five cent coins = 0.05 * x Value in dollars of (x+2) ten cent coins = 0.10 * (x+2) Value in dollars of (28-2x) twenty cent coins = 0.20 * (28-2x) Add all the values, equate it to $3.80 and solve for x.

OpenStudy (anonymous):

it says the answer is 10 but i do not know how to get it

OpenStudy (dumbcow):

you can simplify it by getting rid of decimals (multiply equation by 100) --> 5x + 10(x+2) + 20(28-2x) = 380

OpenStudy (anonymous):

okay i got that #dumbcow but i get 8

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

good job , x=8 read over aum's post and you will see that "x" is number of 5-cent pieces what represents number of 10-cent pieces?

OpenStudy (anonymous):

how/where do i substitute the x=8 @dumbcow

OpenStudy (dumbcow):

number pf 10-cent coins = x+2 "she has two more 10 cent coins than 5 cent coins" we know she has 8 5-cent coins ...